# Steering Large Language Models for Machine Translation Personalization

### v1

@@ 4.2 Steering Baselines
ActAdd is a simple yet effective technique for steering language models. We employ the standard contrastive approach by Rimsky et al. (2024); Scalena et al. (2024) to extract two sets of style-relevant ($\left\{z\right\}^{+}$) and default ($\left\{z\right\}^{-}$) activations from a given model layer using 20 in-context examples demonstrating default behavior (MT) and the desired behavior (H1 or H2 translations), respectively. We then compute the average $\Delta$ steering vector between the two sets of activations, scale it by a factor of $\alpha=2$ which was found effective by previous research (Scalena et al., 2024) and apply it additively to the same model layer during inference.
@@
As an alternative to traditional weight-based Parameter-Efficient Fine-Tuning (PEFT) techniques it learns task-specific interventions applied directly to model activations at inference time (Wu et al., 2024). We apply ReFT to the same personalization-relevant layers identified in Section 3.3 and limit confounding factors by tuning ReFT interventions with the set of 20 examples used for MS prompting.
4.3 Contrastive SAE Steering
Given a set of LLM activations $z_{l}\in\mathbb{R}^{d}$ sourced from the output of layer $l$, where $d$ is the model hidden size, a sparse autoencoder learns a projection to an overcomplete latent space $\mathbb{R}^{m}$, with $m\gg d$, from which a reconstructed version $z^{*}_{l}$ of the original activations is then produced with minimal information loss:
$z^{*}_{l}=\text{SAE}(z_{l})=h(z_{l})W^{l}_{\text{dec}}+b_{\text{dec}}$
[…]
where $W^{l}_{\text{enc}},W^{l}_{\text{dec}}$ are the SAE encoder and decoder modules, and $b_{\text{enc}},b_{\text{dec}}$ are bias values. Our primary interest lies in the sparse latents $h(z_{l})\in\mathbb{R}^{m}$ learned by the SAE encoder, which were empirically found to capture monosemantic and interpretable properties of model inputs.
Contrastive prompt setup
Given a set of paragraphs $\mathcal{D}$ for a novel in the par3 dataset, each instance in it is a tuple:
$\mathcal{D}=\left\{\left<s,\text{{\color[rgb]{0.2734375,0.38671875,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.38671875,0.3828125}H1}},\text{{\color[rgb]{0.51953125,0.73828125,0.61328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.51953125,0.73828125,0.61328125}H2}},\text{{\color[rgb]{0.796875,0.47265625,0.2265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.796875,0.47265625,0.2265625}MT}}_{\text{model}}\right>\right\}$
[…]
capturing respectively the personalized style of interest, and baseline properties of the task. Similarly to the ZS-Exp setup from Section 4.1, we explore two $\mathcal{D}^{-}$ configurations using either $e^{-}=\;$MT (SAE Cont.ht) or $e^{-}=\;$H2  (or H1, if H2 is the personalization target) to assess the effect of baseline choice in steering effectiveness.
@@ 4.3 Contrastive SAE Steering
First, we gather activations $z^{+}_{l}$ and $z^{-}_{l}$ by prompting the model with inputs from the two contrastive sets $\mathcal{D}^{+}$ and $\mathcal{D}^{-}$. Activations are extracted at the last prompt token position from its most informative layer, as identified in Section 3.3. Activations are then converted into sparse latent representations $x^{+}=h(z^{+})$ and $x^{-}=h(z^{-})$, with $x^{+},x^{-}\in\mathbb{R}^{m}$ by the SAE encoder. This procedure is repeated across 20 contrastive examples, resulting in two collections of SAE latent vectors for positive/negative examples:
$\mathcal{X}^{+}$ =$=\left\{x^{+}_{1},x^{+}_{2},\dots,x^{+}_{20}\right\}$
$\mathcal{X}^{-}$ =$=\left\{x^{-}_{1},x^{-}_{2},\dots,x^{-}_{20}\right\}$
@@
To identify discriminative features for personalization in the large set of latents, we employ an information-theoretic approach adapted from Zhao et al. (2025). For each of the inputs, we identify the subset of size $n<m$ including only the SAE active features, i.e. latent dimensions for which the logit is $>0$. We consider logit values in this subset as instances of a random variable $X_{i}\in x$, and calculate the mutual information $I(X_{i},Y)$ between each feature $X_{i}$ and the target binary variable $Y=\left\{+,-\right\}$ corresponding to the style of the provided examples (personalized or non-personalized). A higher $I(X_{i},Y)$ indicates that the $i$-th feature is more informative for discriminating between personalized and default inputs, and can hence be used for steering. A representative sample of 40 latents showing the highest mutual information scores for both personalized ($\left\{X_{i}\right\}^{+}$) and non-personalized ($\left\{X_{i}\right\}^{-}$) examples is selected using this procedure.  [^8] By contrast, traditional SAE-based steering methods only employ features associated with the positive class (Chalnev et al., 2024; Arditi et al., 2024). For every selected latent, we compute its expected logit when personalization is present or absent in provided examples, i.e. $\mathbb{E}^{+}[X_{i}]$ and $\mathbb{E}^{-}[X_{i}]$.
@@
Finally, activations are steered by setting selected latents to their expected value whenever their observed score is below (for the promoted personalized case) or above (for the demoted non-personalized case) the pre-computed average. [^9] Algorithm 1 provides a summary of our SAE-based steering approach. Hence, in the SAE Cont.ht setting we enhance the features relevant to a target personalized style, e.g. $\left\{X_{i}\right\}^{\text{{\color[rgb]{0.2734375,0.38671875,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.38671875,0.3828125}H1}}}$ for H1, and suppress the features $\left\{X_{i}\right\}^{\text{{\color[rgb]{0.796875,0.47265625,0.2265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.796875,0.47265625,0.2265625}MT}}}$, corresponding to the model’s default MT. In SAE Cont.pt, instead, we promote the same H1-related latents while suppressing $\left\{X_{i}\right\}^{\text{{\color[rgb]{0.51953125,0.73828125,0.61328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.51953125,0.73828125,0.61328125}H2}}}$ to steer the model towards H1 personal style. Additionally, we modulate the magnitude of the resulting vector with an $\alpha$ coefficient, which was found to play an essential role in steering effectiveness in previous research (Scalena et al., 2024; Ferrando et al., 2025)
@@ 5.1 Setup
We evaluate our methods on the same three models used for our preliminary evaluation of Section 3. Our selection is guided by the availability of open-source pre-trained SAEs, which can be otherwise computationally expensive to train. For Gemma models, we employ SAEs from the GemmaScope suite (Lieberum et al., 2024); for the Llama 3.1 model we employ the SAE released by (McGrath et al., 2024). GemmaScope SAEs are available for every model layer, enabling us to steer Gemma models on their most informative layers for the task, which we identified in Section 3.3. On the contrary, a single SAE for the 19th layer is available for Llama, hence limiting our evaluation of SAE steering and potentially producing sub-optimal steering result for that model.
@@
H P ${}_{\textsc{flip}}$ [Uncaptioned image] H P ${}_{\textsc{flip}}$ [Uncaptioned image] H P ${}_{\textsc{flip}}$ [Uncaptioned image]
ZS 0.21 0.10 0.05 0.69 0.15 0.08 0.04 0.71 0.24 0.08 0.05 0.70
ZS-Expht 0.30 0.22 0.16 0.68 0.41 0.22 0.18 0.72 0.56 0.23 0.21 0.69
ZS-Exppt 0.20 0.14 0.69 0.23 0.19 0.73 0.30 0.26 0.70
MS 0.37 0.24 0.16 0.69 0.48 0.31 0.27 0.73 0.58 0.32 0.28 0.73
ActAdd 0.27 0.22 0.12 0.67 0.32 0.24 0.20 0.70 0.55 0.36 0.28 0.70
ReFT 0.31 0.22 0.18 0.70 0.46 0.34 0.27 0.67 0.53 0.38 0.26 0.70
SAE Cont.ht 0.39 0.27 0.19 0.70 0.46 0.33 0.29 0.72 0.59 0.31 0.27 0.72
SAE Cont.pt 0.27 0.18 0.69 0.35 0.29 0.73 0.33 0.28 0.72
@@
Metrics
@@
To ensure that our interventions do not result in a degradation of overall translation quality, we also employ [Uncaptioned image] Comet [^10] Unbabel/wmt22-comet-da (Rei et al., 2020) using the personalized translation as reference.
The $\alpha$ trade-off
We begin by verifying the optimal steering intensity $\alpha$ for our SAE steering technique. We primarily focus on results from Gemma 2 2B, for which we ran a comprehensive sweep over all relevant hyperparameters. [^11] Larger models were evaluated using a subset of the best-performing configurations. Details in Appendix C. Figure 3 illustrates the influence of $\alpha$ on MT personalization accuracy and fluency averaged across all translators for all tested languages. For values of $\alpha\leq 3$, performance remains close to that of the MS baseline, indicating that the contrastive method is effectively isolating latents associated with human-like style. As $\alpha$ increases, performance generally exceeds the MS approach, achieving greater control and flexibility in guiding the model’s output with next to no impact of translation quality. However, for $\alpha\geq 10$, we observe a major degradation in [Uncaptioned image] Comet, suggesting an important drop in translation fluency.
Following Ferrando et al. (2025), which also employ SAEs for steering, we experiment with very high alpha values (up to 150), finding the [Uncaptioned image] classifier’s H accuracy approaching 100% for some languages. While this indicates that the contrastive steering is aggressively optimizing toward classifier preferences (Figure 4), the consequent drop in [Uncaptioned image] Comet scores reveals a steep decline in translation quality, often resulting in incoherent or nonsensical generations from a human perspective. [^12] A qualitative evaluation is provided in Appendix D. Ultimately, we identify $\alpha=5$ as an appropriate steering intensity to balance personalization and fluency, and employ it for our main evaluation.
Figure 3: Personalization P and Refer to caption Comet across various steering intensity $\alpha$ for SAE Cont.ht on Gemma 2 2B. The performance of prompting baselines (ZS, MS, Exp) is also reported. Results show a trade-off between steering intensity and translation quality.
Figure 4: Refer to caption Comet and H accuracy across $\alpha$ steering intensity values for Gemma 2 2B, showing a major drop in translation quality for very high intensities ($\alpha\geq 50$).
5.2 Results and Discussion
Table 3 presents performances of tested models across prompting and steering setups, averaged across all languages and personalization targets (H1 and H2 for each language). We find that our SAE Cont.ht and SAE Cont.pt methods generally achieve the best trade-off between personalization accuracy and translation quality, especially for the smaller Gemma 2 2B model. This could be due to the larger models’ superior ability to incorporate in-context information naturally, reducing the relative benefit of explicit steering.
Which contrastive setup is better?
Comparing the two contrastive setups (ht and pt) for the ZS-Exp and SAE Cont. methods, we find that using different human demonstrations as a contrastive baseline in pt generally produces better results for larger models. As for general performance, we conjecture this could be due to the larger models’ improved ability to disentangle personalization-critical factors without explicit guidance. For the smaller Gemma 2 2B, the difference between the two approaches is minimal, suggesting the model cannot fully exploit the examples’ differences.
Do SAE Steering and MS Prompting Impact Activations in a Similar Way?
Since SAE-based approaches perform on par or better than MS, we set out to investigate whether the two methods result in a similar impact on model representations. We collect the modified activations $z_{\text{steer}}$ obtained from the SAE Cont.ht steering setting and evaluate them using the probing classifier trained on MS-conditioned activations we introduced in Section 3.3 for detecting personalization information. Table 4 shows probe accuracy in detecting the positive impact of SAE steering across the three possible outcomes of the steering procedure. We find that the probe corresponding to the SAE layer effectively distinguishes between activations corresponding to successful and unsuccessful SAE steering, despite having been exposed only to MS conditioning during training. This includes both instances for which the classifier prediction is flipped after steering (MT $\rightarrow$ H*), and settings where the conditioning fails (MT $\rightarrow$ MT). In settings where the original output already matches human style (H* $\rightarrow$ H*), the probe obtains lower accuracy with broader confidence intervals, denoting higher uncertainty. These findings suggest that the SAE’s latents we extract through our contrastive method are meaningfully connected to the stylistic patterns embedded in the multi-shot examples, providing evidence that our intervention influences the internal representations of the model, aligning them to the natural effect of the MS approach.
MT $\rightarrow$ H* MT $\rightarrow$ MT H* $\rightarrow$ H*
@@ 5.2 Results and Discussion
±$\pm 0.19$
Table 4: Probing accuracy on SAE Cont.ht-steered activations averaged across languages and H1/H2 translators. Probes trained on MS activations reliably detect the impact of SAE Cont. steering over model generations, suggesting similar mechanisms.
@@ 6 Conclusion and future work
We conducted a broad evaluation of various prompting and steering approaches for personalizing LLM-generated translations. Our evaluation targets a practical, real-world application of literary translation and addresses the underexplored challenge of steering LLM generations in a linguistically rich and stylistically sensitive domain. Through comprehensive evaluation across multiple languages, novels, and models, we demonstrate that our proposed SAE-based approach outperforms prompting and alternative steering techniques.
Although faithfully replicating individual human translation styles remains a highly challenging task, our approach achieves strong alignment with human translation quality, as reflected in both general human-likeness and translator-specific personalization metrics. These results highlight the method’s robustness and its potential to support high-fidelity translation workflows in real-world settings. Concretely, these results have important implications in the development of personalized MT systems based on LLMs. In particular, the notable effectiveness of our proposed approach on smaller models might enable MT personalization in lower-resource settings, easing further research on how personalization information is encoded and produced by language models.
Future work will focus on improving the interpretability of the learned SAE latents, to uncover their learned monosemantic features. Additionally, we plan to explore the integration of our method with larger language models, where increased capacity may further enhance the precision and fluency of personalized translations.
@@
While our work demonstrates the potential of steering LLMs for MT personalization using sparse autoencoders, we acknowledge several limitations.
Firstly, our findings’ generalizability is constrained by the scope of our experiments. We focused on literary translation into English from seven specific source languages and evaluated three LLMs of relatively small size. Consequently, the observed effectiveness of SAE-based steering and the identified optimal layers for intervention may not directly transfer to other language pairs, significantly different model architectures or sizes, or distinct domains beyond literary texts. Further research is needed to assess the robustness of our approach across a broader range of linguistic and modeling contexts.
Secondly, the computational overhead associated with sparse autoencoders presents a practical challenge. Although we utilized pre-trained SAEs in our study, the initial training of these components is resource-intensive. This could limit the accessibility and scalability of our proposed method, particularly for researchers or practitioners with limited computational resources or when frequent retraining for new models or tasks is required. The current availability of pre-trained SAEs also restricts model choice, as seen with the Llama 3.1 8B model where an SAE was only available for a potentially sub-optimal layer.
Finally, our investigation primarily focused on downstream performance and the impact of various personalization strategies on model representations. However, we did not pursue a mechanistic understanding of the "personalization circuits" within the LLMs. Future work could adopt a more fine-grained, mechanistic interpretability approach to study how specific SAE latents or combinations thereof encode and manipulate nuanced stylistic features, thereby providing deeper insights into the underlying processes of LLM personalization.
Appendix A List of novels used
To ensure a diverse and representative evaluation, we select novels spanning a variety of linguistic families and cultural backgrounds. Our dataset includes Romance languages such as Italian (Pinocchio) and French (Around the World in Eighty Days), as well as Germanic languages like Dutch (The Diary of a Young Girl) and German (Beware of Pity). To evaluate our setup on non-Latin scripts and distinct linguistic structures, we also include Russian (Crime and Punishment), Japanese (No Longer Human), and Chinese (Dream of the Red Chamber). Table 5 summarizes the number of paragraphs employed in the evaluation of each language.
Lang Novel name Train Val Test ICL
IT Pinocchio 745 82 107 20
FR Around the World in Eighty Days 829 92 120 20
NL The Diary of a Young Girl 769 85 110 20
DE Beware of Pity 606 67 96 20
RU Crime and Punishment 1517 168 224 20
JA No Longer Human 652 40 81 20
ZH Dream of the Red Chamber 694 76 92 20
Table 5: Each number corresponds to a single instance in the dataset. When the dataset is used for training, each instance is associated with three distinct labels - H1, H2, and MT - thus the total number must be multiplied by three.
Appendix B Experiments reproducibility

### v2

@@ 4.2 Steering Baselines
ActAdd is a simple yet effective technique for steering language models. We employ the standard formulation by Rimsky et al. (2024), contrasting a set of positive activations $\left\{z\right\}^{+}$, here the style-aligned $H^{+}$ translations, with negative activations $\left\{z\right\}^{-}$, here the default MT output. We perform steering on the style-relevant layers identified in Section˜3.3 by computing the average $\Delta$ steering vector as the difference-in-means for the two sets of activations using 20 in-context examples, and applying it additively to the same model layer during inference. We use a factor $\alpha=2$ to control steering intensity, as this was found effective in previous research (Scalena et al., 2024).
@@
ReFT was recently introduced as an alternative to traditional weight-based Parameter-Efficient Fine-Tuning (PEFT) techniques, aiming to learn task-specific interventions applied directly to model activations at inference time (Wu et al., 2024). As for ActAdd, we apply ReFT to the style-relevant layers identified in Section˜3.3 and limit confounding factors by tuning ReFT interventions with the same set of 20 examples used for MS prompting.
4.3 Contrastive SAE Steering (SAE Cont.)
We now introduce our proposed contrastive formulation for steering translation style with SAE latents. Given the set of paragraphs $\mathcal{D}$ for a novel in the par3 dataset, each instance in it is a tuple $\left<s,H^{+},H^{-},\text{{\color[rgb]{0.796875,0.47265625,0.2265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.796875,0.47265625,0.2265625}MT}}\right$>, where $s$ is the non-English source sentence. Similar to previous methods, we employ a contrastive approach to extract SAE latents that are most active in the presence of the selected translator’s style, while simultaneously controlling for generic features occurring in natural model outputs.
@@
First, we gather activations $z^{+}_{l}$ and $z^{-}_{l}$ by prompting the model with tuples $\left<s,H^{+}\right$>, $\left<s,\text{{\color[rgb]{0.796875,0.47265625,0.2265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.796875,0.47265625,0.2265625}MT}}\right$>, respectively. Activations are extracted at the last prompt token position from the most style-related layer, as identified in Section 3.3. Activations are then converted into sparse feature dictionaries $x^{+}=h(z^{+})$ and $x^{-}=h(z^{-})$, with $x^{+},x^{-}\in\mathbb{R}^{m}$ and $h$ representing the the SAE encoder module. This procedure is repeated separately across 20 contrastive examples, resulting in two collections of SAE latent vectors for positive/negative examples, $\mathcal{X}^{+}=\left\{x^{+}_{1},x^{+}_{2},\dots,x^{+}_{20}\right\}$ and $\mathcal{X}^{-}=\left\{x^{-}_{1},x^{-}_{2},\dots,x^{-}_{20}\right\}$
@@
To identify discriminative features for personalization in the large set of latents, we adapt the information-theoretic approach proposed by Zhao et al. (2025). For each of the inputs, we identify the subset of size $n\ll m$ of SAE active features, i.e. latent dimensions for which the logit is >0>0. We consider logit values in this subset as instances of a random variable $X_{i}\in x$, and calculate the mutual information $I(X_{i},Y)$ between each feature $X_{i}$ and the target binary variable $Y=\left\{+,-\right\}$, representing whether the text matches the desired $H^{+}$ style or not. A higher $I(X_{i},Y)$ indicates that the $i$-th feature is more informative for discriminating between personalized and default inputs, and can hence be used for steering. A representative sample of 40 latents showing the highest mutual information scores for both personalized ($\left\{X_{i}\right\}^{+}$) and non-personalized ($\left\{X_{i}\right\}^{-}$) examples is selected using this procedure. This procedure differs from traditional SAE-based steering methods, which employ only features associated with the positive class (Chalnev et al., 2024; Arditi et al., 2024), in that it encourages the selection of discriminative features, both positive and negative. For every selected latent, we use its expected logit when personalization is present or absent, given the set of provided examples, i.e. $\mathbb{E}^{+}[X_{i}]$ and $\mathbb{E}^{-}[X_{i}]$.
@@
Finally, activations are steered by clipping selected latents’ values to their expected value, hence promoting $H^{+}$-aligned latents $\left\{X_{i}\right\}^{H^{+}}$ while demoting MT-aligned ones $\left\{X_{i}\right\}^{\text{{\color[rgb]{0.796875,0.47265625,0.2265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.796875,0.47265625,0.2265625}MT}}}$. [^7] Algorithm 1 summarizes our SAE steering approach. Additionally, we modulate the magnitude of the resulting vector with an $\alpha$ coefficient, which was found to play an essential role in steering effectiveness in previous research (Scalena et al., 2024; Ferrando et al., 2025)
@@ 5.1 Setup
We evaluate our methods on the Llama 3.1 and Gemma models previously used in Section˜3. Our selection is guided by the availability of open-source pre-trained SAEs, which can be otherwise computationally expensive to train. For Gemma models, we employ SAEs from the GemmaScope suite (Lieberum et al., 2024), while for the Llama 3.1 model we use the SAE by McGrath et al. (2024). GemmaScope SAEs are available for every model layer, enabling us to steer Gemma models on their most informative layers for the task, which we identified in Section˜3.3. On the contrary, a single SAE for the 19th layer is available for Llama, hence limiting our evaluation of SAE steering and potentially producing sub-optimal steering results for that model.
@@
H P $_{\textsc{flip}}$ [Uncaptioned image] tok/s H P $_{\textsc{flip}}$ [Uncaptioned image] tok/s H P $_{\textsc{flip}}$ [Uncaptioned image] tok/s
ZS 0.21 0.10 0.05 0.69 39.8 0.15 0.08 0.04 0.71 25.7 0.24 0.08 0.05 0.70 25.9
ZS-Exp. 0.30 0.22 0.16 0.68 41.3 0.41 0.22 0.18 0.72 24.6 0.56 0.23 0.21 0.69 25.5
MS 0.37 0.24 0.16 0.69 36.0 0.48 0.31 0.27 0.73 16.8 0.58 0.32 0.28 0.73 17.2
ActAdd 0.27 0.22 0.12 0.67 40.2 0.32 0.24 0.20 0.70 25.3 0.55 0.36 0.28 0.70 24.2
ReFT 0.31 0.22 0.18 0.70 40.7 0.46 0.34 0.27 0.67 25.5 0.53 0.38 0.26 0.70 24.7
SAE Cont. 0.39 0.27 0.19 0.70 41.1 0.46 0.33 0.29 0.72 25.1 0.59 0.31 0.27 0.72 23.1
Table 4: Averaged metric scores across all tested languages (per-language breakdown in Appendix˜B). H: human style accuracy, i.e. p(p(H1)+p()+p(H2)). P: personalization accuracy $p(H^{+})$ for the desired style. $_{\textsc{flip}}$: % of segments for which style conditioning flips the classifier prediction from MT to $H^{+}$. $\alpha=5$ is used for SAE Cont. results.
Metrics
We evaluate our approaches on a held-out test set sourced from the par3 dataset. For assessing style accuracy, we use the [Uncaptioned image] classifiers described in Section˜3.1. We define three submetrics employing the classifier probability distribution over the three classes (MT, H1, H2) to better analyze different aspects of model outputs. First, we define the H accuracy as the [Uncaptioned image] classifier’s total probability assigned to human-like translations, p(p(H1)+p()~+~p(H2)), thereby measuring generic human-like features of the text. We instead define P accuracy as the probability of the desired $H^{+}$ class, which corresponds to specific stylistic traits of a human translator. Finally, we use $_{\textsc{flip}}$ to indicate the proportion of examples for which the [Uncaptioned image] classifier prediction flips from MT  to $H^{+}$ after prompting or steering is applied, pinpointing examples for which the method has an observable effect on generation style. To ensure that our interventions do not result in a degradation of overall translation quality, we also employ [Uncaptioned image]  Comet [^8] Unbabel/wmt22-comet-da (Rei et al., 2020) using the $H^{+}$ as reference.
Figure 3: Effect of various steering intensity $\alpha$ on style accuracy and translation quality for Gemma 2 2B. Top: P accuracy for SAE Cont. and prompting baselines (MS, ZS-Exp and ZS). Bottom: H accuracy for high $\alpha$ showing a steep drop in translation quality while style accuracy increases.
The $\alpha$ trade-off
We begin by verifying the optimal steering intensity $\alpha$ for our SAE steering technique. We primarily focus on results from Gemma 2 2B, for which we ran a comprehensive sweep over all relevant hyperparameters. [^9] Larger models were evaluated using a subset of the best-performing configurations. Details in Appendix B. Figure 3 (top) shows the influence of $\alpha$ on MT personalization accuracy and fluency, averaged across all translators and all tested languages. For values of $\alpha\leq 3$, performance remains close to that of the MS baseline, indicating that the contrastive method is effectively isolating latents associated with the desired style. As $\alpha$ increases, performance generally exceeds the MS approach, achieving greater stylistic accuracy with minimal impact on translation quality. However, for $\alpha\geq 10$, we observe a degradation in translation fluency, represented by [Uncaptioned image]  Comet.
Following the SAE steering setup of Ferrando et al. (2025), we also experiment with very high $\alpha$ (up to 150), obtaining H accuracies approaching 100% for some languages but also resulting in abysmal translation quality, with generally nonsensical outputs (Figure˜3 bottom). These results suggest that SAE features identified by our contrastive approach correspond to human-like constructions, provided that their upweighting raises the [Uncaptioned image] classifier confidence in predicting H1 or H2, even when the resulting quality is clearly lacking. [^10] A qualitative evaluation is provided in Appendix C. We leave an exploration on using such examples to further improve style classifiers to future work. Ultimately, we identify $\alpha=5$ as an appropriate steering intensity that balances personalization and fluency, and employ it in our main evaluation.
5.2 Results and Discussion
Table 4 presents the performance of tested models across prompting and steering setups, averaged across all languages and styles, i.e., using H1 and H2 as $H^{+}$ for each language. We find that contrastive SAE steering (SAE Cont.) yields a good balance between style accuracy and translation quality, with results comparable to the strong MS prompting baseline. Notably, SAE Cont. outperforms all other tested methods for the smaller Gemma 2 2B model. We conjecture this could be due to the larger models’ ability to incorporate in-context information more effectively, reducing the benefits of ad-hoc interventions.
Is steering computationally efficient?
@@ 5.2 Results and Discussion
Figure 4: Personalization accuracy P (top) and inference speed (Tokens/s) (bottom) across in-context demonstration counts, using Gemma-2 9B for Russian $\rightarrow$ English translation. More results in Appendix˜E.
Is SAE steering an “internalized” MS prompting?
@@
MT $\rightarrow H$ MT $\rightarrow$ MT $H\rightarrow H$
@@
$\pm 0.19$
Table 5: Estimated $p(H)$ for SAE-steered activations using Section˜3.3 binary probes trained on MS-conditioned activations, across three settings identified by [Uncaptioned image] classifier input/output labels.
@@ 6 Conclusion and future work
We conducted a broad evaluation of various prompting and steering approaches for personalizing LLM-generated translations. Our evaluation targets a practical, real-world application of literary translation, addressing the underexplored challenge of steering LLM generations in a linguistically rich and stylistically sensitive domain. Through comprehensive evaluation across multiple languages, novels, and models, we demonstrate that our proposed SAE-based approach outperforms prompting and alternative steering techniques, resulting in a good balance of style conditioning and translation quality at negligible inference-time cost. In particular, our proposed method scales effectively to large amounts of examples, bypassing the context limitations and slowdowns of in-context demonstrations.
Although faithfully replicating individual human translation styles remains a highly challenging task, our approach achieves strong alignment with human translation quality and translator-specific style features. In particular, the effectiveness of our proposed approach on smaller models might enable MT personalization in smaller-scale computational settings, easing further research on how style information is encoded and produced by language models.
@@
While our work demonstrates the potential of steering LLMs for MT personalization using sparse autoencoders, we acknowledge some limitations in our setup.
Firstly, the generalizability of our findings is constrained by the scope of our experiments. We focused on literary translation from seven specific source languages into English and evaluated three LLMs of relatively small size. Consequently, the effectiveness of SAE-based steering and the identified optimal layers for intervention may not directly transfer to other language pairs, significantly different model architectures or sizes, or distinct domains beyond literary texts. Further research is needed to assess the robustness of our approach across a broader range of linguistic and modeling contexts.
Secondly, the computational overhead associated with training sparse autoencoders presents a practical limitation to the widespread adoption of our proposed method. Although many open-source pre-trained SAEs are now available, training new SAEs from scratch may be required for newer models, which might pose a problem for researchers with limited computational resources. The current availability of pre-trained SAEs also restricts model choice, as seen with the Llama 3.1 8B model where the SAE we selected was available for a single, potentially suboptimal layer.
Finally, our investigation primarily focused on downstream performance and the impact of various personalization strategies on model activations. However, we did not pursue a mechanistic understanding of the "personalization circuits" within the LLMs. Future work could adopt a more fine-grained, mechanistic interpretability approach to study how specific SAE latents or combinations thereof encode and manipulate nuanced stylistic features, thereby providing deeper insights into the underlying processes of LLM personalization.
Appendix A Experimental Reproducibility